The sum of three numbers is 23. If the third integer is doubled the sum is 35. The second is 3 or more than the first integer. Find the three simultaneous equations and the three integers.
Answer by richwmiller(17219)

y=x+3
x+y+z=23
x+y+2z=35
x+y+12=23
z=12
x+y=11
x+x+3=11
2x=8
x=4
y=7